OpenStudy (anonymous):

state how many complex and real zeros the function has f(x)=x^3+5x^2+x+5

Parth (parthkohli):

\[x^3+ 5x^2 + x + 5\]\[= x^2( x + 5) + 1(x + 5) = (x^2 + 1)(x + 5)\]